OpenStudy (anonymous):

I have a total of 13 dimes and quarters which equals $2.05. How many quarters and dimes do I have? I don't know now to do this, please help!

OpenStudy (supernaturallvr24):

5 quarters and 8 dimes.

OpenStudy (anonymous):

How did you get that though?

OpenStudy (chealyn98):

D + g =13 0.10d + 0.25q = 2.05 ( number of dimes and quarters) ( their value) - 0.10d - 0.10q = -1.3 0.10d + 0.25q = 2.05 0.15/0.15q 0.75/0.10d Q= 5 D= 8